OpenStudy (anonymous):

PLEASE HELP ME IM STUCK ON THIS PRE CALC QUESTION??? PLEASE HELP ME FINISH MY WORK! Find an equation for the nth term of the arithmetic sequence. -20, -16, -12, -8, ... Here is what i have so far: d=4 a1=-20 c=a1-d c=-24 an=dn+c an=4n-24 What are my next steps? Please help!

OpenStudy (anonymous):

These are my answer choices A. an = -20 x 4(n - 1) B. an = -20 + 4(n - 1) C. an = -20 + 4(n + 2) D. an = -20 + 4(n + 1)

jimthompson5910 (jim_thompson5910):

\[\Large \text{nth term} = (\text{first term}) + (\text{common difference})(n-1)\] \[\Large a_n = a_1 + d(n-1)\]

OpenStudy (anonymous):

So, an= -20 + 4(n-1)?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

What is that formula called? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm not sure of the specific name, but you can consider it the nth term of the arithmetic sequence formula

OpenStudy (anonymous):

Alright, thank you very much. I did not have that in my notes. Thanks! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome